Water enters the heater at 30oC and leaves at 150oF, what is the temperature difference in oC? answer in 2 decimal places

1
Expert's answer
2021-08-16T14:44:35-0400

"T_1 = 30 \;°C \\
T_2 = 150 \;°F \\
T_2 = (150 \;°F -32) \times \frac{5}{9}=65.55 \;°C \\
ΔT= 65.55 -30 = 35.55 \;°C"